OpenStudy (anonymous):

A company builds Electronic Entertainment Centers. It costs $8,700 to build 10 Electronic Entertainment Centers and $14,300 to build 20 Electronic Entertainment Centers. Which equation models the cost, C(x), as a linear function of the number of Electronic Entertainment Centers, x? Answer C(x) = 3100x + 560 C(x) = 560x + 3100 C(x) = 560x - 3100 C(x) = 3100x - 560

OpenStudy (unklerhaukus):

\[C(10)=$8,700\]\[C(20)=$14,300\]

OpenStudy (anonymous):

then what?

OpenStudy (unklerhaukus):

the first option is this \[C(x) = 3100x + 560\] if we try x=10 we get \[C(10)=3100\times10+560=$31,560\ne$8,700\] so that option is not correct

OpenStudy (anonymous):

alright

OpenStudy (unklerhaukus):

try another option , \[\cancel{C(x) = 3100x + 560}\] \[C(x) = 560x + 3100\] \[C(x) = 560x - 3100 \] \[C(x) = 3100x - 560\]

OpenStudy (unklerhaukus):

plug in x=10 and try to get 8700 if that works you still have to check that x=20 gives 14,300 as-well, since the equations are straight lines two points are enough to define the line

OpenStudy (anonymous):

so which 1 is it?

OpenStudy (unklerhaukus):

can you see how i checked the first equation for c(x)?

OpenStudy (anonymous):

somewhat

OpenStudy (unklerhaukus):

i tested the point in the equations with an input and checked if the out put was what i wanted

OpenStudy (anonymous):

ok

OpenStudy (unklerhaukus):

to test the second function \(C(x)=560x+3100\) \[C(10)=560\times10+3100=? \] does this equal $8,700?

OpenStudy (anonymous):

yes

OpenStudy (unklerhaukus):

so one point on the line \(C(x)=560x+3100\) is the same is the other point the same also? is \(C(20)=14300\)?

OpenStudy (anonymous):

yea

OpenStudy (unklerhaukus):

so you have prove two point on the line are right, now since both lines are linear , they must be the same line.

OpenStudy (anonymous):

alright thanks

OpenStudy (unklerhaukus):

so B.